# MARM: Unlocking the Future of Recommendation Systems through Memory Augmentation and Scalable Complexity

### v2

@@ Abstract.
Scaling-law has guided the language model design for past years, to estimate expected model performance with respect to the size of learnable parameters and training samples data-scales, e.g. GPTs. It is worth noting that the scaling laws of NLP cannot be directly applied to recommendation systems due to the following reasons: (1) The amount of training samples and model parameters is typically not the bottleneck for the model. Our recommendation system can generate over 50 billion user samples daily, and such a massive amount of training data can easily allow our model parameters to exceed 200 billion, surpassing many LLMs (about 100B). (2) It is essential to control computational complexity FLOPs carefully in recommendation system. In training, we need to process a vast number of recommendation samples every day. During online inference, we must respond within milliseconds (LLMs take usually a few seconds). Considering the above differences with LLM, we can conclude that: for a RecSys model, compared to model parameters, the computational complexity FLOPs is a more expensive factor that requires careful control.
In this paper, we propose our milestone work, MARM (Memory Augmented Recommendation Model), which explores a new cache scaling-laws successfully. By caching part of complex module calculation results, Our MARM extends the single-layer attention-based sequences interests modeling module to a multiple-layer setting with minor inference complexity FLOPs cost (i.e, module time complexity $\mathcal{O}(n^{2}*d)\rightarrow\mathcal{O}(n*d)$). Equipped with the cache idea, our MARM solution significantly overcomes computational bottlenecks, and could seamlessly empower all user sequences interest extracting modules, and even other models. To support our MARM, we construct a 60TB cache storage center for offline training and online serving. Comprehensive experiment results show that our MARM brings offline 0.43% GAUC improvements and online 2.079% play-time per user gains. Our MARM has been deployed on Kwai short-videos platform, serving tens of millions of users daily.
Scaling-Laws; Cache Technique; Low-Resource Large Model;
ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. (a) A toy example of ranking model. (b) Performance gap in Streaming Data Training. Warming up from the online base model, a 50% sample downsampling leads to a continuous decline in model performance. On the other hand, there is a significant performance gap between a cold-start model trained from scratch and the online base model, but the gap decreases as the training data increases. (c) Parameter distributions. The parameters of an LLM model (blue) mainly are located in ‘dense-activated’ transformer-based DNN module, while ranking models (green) have a parameter distribution primarily composed of ‘sparse-activated’ feature embeddings.
@@
In recent years, the scaling of model size and data-scale has been a vital point in various fields, e.g., NLP, CV, and RecSys. According to OpenAI Scaling-laws technique report(Kaplan et al., 2020), the volume of data and the breadth and depth of models increase, the performance of the models follows a certain power-law improvement. Following the excited Scaling-laws tenet, many Transformer-based large models are proposed and achieve remarkable performance, e.g., ChatGPT for conversation, Qwen-VL(Bai et al., 2023) for multi-modal understanding, DeepSeek-Coder(Guo et al., 2024) for code generation, and Kling for Video generation. There are also some efforts to valid Scaling-laws in RecSys area (e.g., Wukong(Zhang et al., 2024a), HSTU(Zhai et al., 2024)), but these methods make strong assumptions on features engineering or model architectures, e.g., remove all statics features, replace entire model modules by Transformer. Actually, the common wisdom always formed RecSys ranking model designing as follows: first crafting hundreds elaborated features as model input and then use multi-task prediction module to obtain the click/like/comment/others scores for a user and item candidate pair (in Figure 1(a)). Nevertheless, the recent RecSys Scaling-laws studies are greatly changing such learning paradigms, which is difficult to directly deploy them in real online RecSys ranking models. In this paper, we focus on exploring the Scaling-laws under the above wide-used ranking model architecture. Due to differences in ranking model architecture and usage with pure Transformer-based LLMs, we conclude three elements that the reasons why NLP Scaling-laws do not match to RecSys ranking model:
@@
  • For the training samples: different from NLP models having a stable data corpus to train an LLM from scratch, the industrial RecSys model always follows a streaming training paradigm. At Kuaishou, our model is streaming trained over 50 billion user logs daily, here we give three model variants training details, e.g., the ‘online base model’, ‘online base model with 50% training data downsampling’, and ‘cold-start’, where the ‘online base model’ is trained over years. From Figure 1(b), we can draw conclusions that: (1) training from scratch will hurt performance significantly; (2) although the model parameters is warm-start from a several years well-trained model, down-sampling the real-time training samples will still have performance degradation. Such phenomenons validate the RecSys model is data-hunger for infinite data to capture real-time users’ preferences. Therefore, the analysis of training data-scale is unnecessary in industry.
@@
  • For the learning parameters: As shown in Figure 1(c), instead of LLM parameters are mainly located in ‘dense-activated’ Transformer-based DNN module (about 100 Billion) but with a small ‘sparse-activated’ word token parameters (¡0.01B). Ranking models always have contrary parameter distribution that: the major learnable parameters are concentrated on the ‘sparse-activated’ feature parameters (¿200B) while ‘dense-activated’ DNN parameters space (about 0.1B) is much smaller. Where the ‘sparse’ means only small part parameters will activate to involve calculating (i.e., we only need to look up few words tokens for a sentence in LLM; we only need to look up one user/item ID for a user-item training sample in RecSys), and the ‘dense’ means the calculation flow parameters that will fully activate to estimate final results. Indeed, in terms of the amount of learnable parameters, our ranking model (¿ 200B) far exceeds many LLMs (around 100B), which indicates that learnable parameters are not the bottleneck of ranking models.
@@
  • For the inference complexity, FLOPs: Actually, the RecSys model inference complexity FLOPs is much smaller than LLM and has its upper bound. This is because our model needs to process hundreds of millions of requests efficiently, and that’s why our dense-activate DNN module is much smaller (about 0.1B). Hereby we could not blindly increase the inference complexity, to ensure our service’s stability and robustness of processing each request in milliseconds (LLMs are usually a few seconds).
@@
Based on the analysis of the above three aspects, compared to LLM, we can observe that our ranking model has its advantages and disadvantage are: (1) ‘unlimited’ streaming data, (2) massive parameter storage, (3) relatively low-FLOPs DNN module. In other words, for ranking model, the data and storage resources are cheap relatively, but inference computing resources are cautiously expensive. Motivates by such point, we consider that can we use the model the strengths in data and storage bootstrapping its weakness in computing. In other words, can we cache part of complex module calculation results to degenerate its time complexity?
To answer the question, we propose our milestone work, MARM (Memory Augmented Recommendation Model), which achieves a new RecSys scaling-laws between the cache size and model performance. At MARM, we extend one of the most important user interests extraction modules in industry ranking model, which aims to calculate users historical item importance with the candidate item, as shown at the bottom of Figure 1(a). To the best of our knowledge, in implementing this module, many outstanding methods (e.g., DIN(Zhou et al., 2018), SIM(Pi et al., 2020), SDIM(Cao et al., 2022), TWIN(Chang et al., 2023)) are utilized a single-layer target-attention mechanism, as shown in Figure 2(a). Intuitively, such module could be further enhanced by stacking multi-layer self-attention before the final target-attention, as shown in Figure 2(b). Unfortunately, the self-attention mechanism time complexity $\mathcal{O}(n^{2}*d)$ is much higher than target-attention mechanism time complexity $\mathcal{O}(n*d)$ ($n$ denotes the sequence length, $d$ denotes representation dimension), which is more sensitive on longer sequences (e.g., $n>1000$) and will result in a large increasing in FLOPs. In online-serving, our ranking model needs to predict 50$\sim$8000 user/item-candidate pairs at same time to find the best dozens items for a user, causing the naive multi-layer attention mechanism will produce a lot of repeated calculations, which will bring heavy pressure for our system (in Figure 2(c)).
@@
If we cache the frequently used masked self attention module results, the complex masked self attention layer can be replaced by the simple target attention layer to efficiently predict different item candidates well, as shown in Figure 2(d). In this way, our MARM extends the single-layer attention-based sequences interests modeling module to a multiple-layer setting with minor inference complexity FLOPs cost. Specifically, our MARM solution significantly overcomes computational bottlenecks, and could seamlessly empower all user sequences interest extracting modules. Moreover, based on the MARM cache results, we find it can contribute to other models, such as retrieval and cascading models. In our experiments, we fully explore the scaling-laws between cache size and model performance, seeking a perfect balance in terms of cached attention depth, sequence length and embedding dimension.
The main contributions of this work are:
@@
  • We are the first work to explore new scaling-laws under the wide-used ranking model architecture from a fresh cache-performance perspective.
@@
  • We devise a simple-yet-efficient approach, MARM, which utilizes the strengths of ranking model in data and storage to degenerate complex masked-self-attention time complexity that we successfully extend the single-layer attention-based sequences interests modeling module to a multiple-layer setting with minor inference complexity FLOPs cost.
@@ 2. Methodology
In this section, we do not explain the total details of our ranking model’s architectures or losses, its sketch illustration is shown in Figure 1(a). Specifically, for better understanding, we dive into the MARM module and explain how it works only. We first introduce the base MARM workflow, then the MARM with SIM extensions, and finally show MARM how to improve other models. Further, we also discuss the cache scaling-law and management strategies, seeking a balance between model complexity and performance.
@@ 2.1. MARM Workflow
This section describes MARM workflow under the latest exposure item sequence, to build a multi-layer decoder-only Transformer architecture to capture user interests with respect to a certain target item. Actually, our MARM mainly consists of four parts:
@@
    An external cache memory storage to look up or save results.
@@
    Sending the intermediate results to cache store.
@@ 2.1.1. Sequence Generator
[…]
@@ 2.1.2. Cache Memory Storage Look-Up
[…]
@@
    where $\mathbf{IidY}\in\mathbb{R}^{F}$, $F$ is a hyper-parameter to control the feature dimension. The $[\mathbf{Iid1},\dots,\mathbf{Iidn}]$ is the the bottom target-attention input in Figure 2(a). It is worth noting that viewed items could fulfill some other attributes, such as author ID, tags, and user interaction labels. In this paper, for simplicity, we will only retain the item ID to represent them.
@@
    Un-Learnable Cached results: Considering that we have cached $L$-layer intermediate results in an external key-value memory storage, here we need to look up those results precisely. Therefore, we devise a hash strategy to generate ‘key’ to reflect the meaning: cached value of user-item pair at layer $i$. for example, given user UidX, and item sequence [Iid1,…] and layer depth $i$, we could generate the hash key list as:
@@
[…]
@@
    where $\mathbf{Iid1}^{i}_{\texttt{UidX}}\in\mathbb{R}^{d}$, and $d$ is attention module dimension. As a result, combining all layers’ cache results we have:
@@
Figure 3. MARM realtime cache update&use workflow;The blue box illustrates the process of the first displayed item being updated into the cache, followed by the subsequent retrieval and usage from the cache.

### v3

# MARM: Unlocking the Recommendation Cache Scaling-Law through Memory Augmentation and Scalable Complexity

@@ Abstract.
Scaling-law has guided the language model design for past years, e.g., GPTs, enabling the estimation of expected model performance with respect to the size of learnable parameters and the scale of training samples. It is worth noting that the scaling laws of NLP cannot be directly applied to recommendation systems due to the following reasons: (1) The amount of training samples and model parameters is typically not the bottleneck for the model. Our recommendation system can generate over 50 billion user samples daily, and such a massive amount of training data can easily allow our model parameters to exceed 200 billion, surpassing many LLMs (about 100B). (2) It is essential to control FLOPs carefully in recommendation system. In training, we need to process a vast number of recommendation samples every day. During online inference, we must respond within milliseconds (LLMs usually take a few seconds). Considering the above differences with LLM, we can conclude that: for a RecSys model, compared to model parameters, the FLOPs is a more expensive factor that requires careful control.
In this paper, we propose our milestone work, MARM (Memory Augmented Recommendation Model), which explores a new cache scaling-law successfully. By caching part of complex module calculation results, our MARM extends the single-layer attention-based sequences interests modeling module to a multi-layer setting with minor inference FLOPs cost (i.e, module time complexity $\mathcal{O}(n^{2}*d)\rightarrow\mathcal{O}(n*d)$). Equipped with the cache idea, our MARM solution significantly overcomes computational bottlenecks and can seamlessly empower all interest extraction modules for user sequences, and even other models. To support our MARM, we construct a 60TB cache storage center for offline training and online serving. Comprehensive experiment results show that our MARM brings offline 0.43% GAUC improvements and online 2.079% play-time per user gains. Our MARM has been deployed on a real-world short-video platform, serving tens of millions of users daily.
journalyear: 2025copyright: ccconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761015isbn: 979-8-4007-2040-6/2025/11ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. (a) A toy example of ranking model. (b) Performance gap in Streaming Data Training. Warming up from the online base model, a 50% sample downsampling leads to a continuous decline in model performance. (c) Parameter distributions. The parameters of an LLM model (blue) are mainly located in ‘dense-activated’ transformer-based DNN module, while ranking models (green) have a parameter distribution primarily composed of ‘sparse-activated’ features.
@@
In recent years, the scaling of model size and data-scale has been a vital point in various fields, e.g., NLP, CV, and RecSys. According to OpenAI Scaling-laws technique report(Kaplan et al., 2020), as the volume of data and the breadth and depth of models increase, the performance of the models follows a certain power-law improvement. Following the excited Scaling-laws tenet, many Transformer-based large models are proposed and achieve remarkable performance, e.g., ChatGPT for conversation, Qwen-VL(Bai et al., 2023) for multi-modal understanding, DeepSeek-Coder(Guo et al., 2024) for code generation, and so on. There are also some efforts to validate Scaling-laws in RecSys area (e.g., Wukong(Zhang et al., 2024a), HSTU(Zhai et al., 2024)), but these methods make strong assumptions on feature engineering or model architectures, such as removing all static features or replacing entire model modules with Transformers. Actually, the common wisdom always formed RecSys ranking model designing as follows: first crafting hundreds of elaborated features as model input and then using a multi-task prediction module to obtain the click/like/comment/others scores for a user and item candidate pair (in Figure 1(a)). Nevertheless, recent RecSys Scaling-laws studies are greatly changing such learning paradigms, making it difficult to deploy them in real online RecSys ranking models directly. In this paper, we focus on exploring the Scaling-laws under the above widely-used ranking model architecture. Due to differences in ranking model architecture and usage with pure Transformer-based LLMs, we conclude three reasons why NLP Scaling-laws do not match to RecSys ranking model:
@@
  • For the training samples: different from NLP models having a stable data corpus to train LLM from scratch, the industrial RecSys model always follows a streaming training paradigm. At real-world RecSys, our model is streaming trained over 50 billion user logs daily. Here we give three model variants training details: the ‘online base model’, ‘online base model with 50% training data downsampling’, and ‘cold-start’, where the ‘online base model’ is trained over years. From Figure 1(b), we can draw conclusions that: (1) Training from scratch will hurt performance significantly; (2) although the model parameters are warm-started from a several-year well-trained model, down-sampling the real-time training samples will still have performance degradation. Such phenomena validate that the RecSys model is data-hungry for infinite data to capture real-time users’ preferences. Therefore, the analysis of training data-scale is unnecessary in industry.
@@
  • For the learning parameters: As shown in Figure 1(c), LLM parameters are mainly located in ‘dense-activated’ Transformer-based DNN module (about 100 Billion), while the ‘sparse-activated’ word token parameters are small (¡0.01B). Ranking models always have opposite parameter distribution: major learnable parameters concentrate on ‘sparse-activated’ feature parameters (¿200B) while ‘dense-activated’ DNN parameters (about 0.1B) are much smaller. Where ‘sparse’ means only small part parameters will activate to involve calculating (i.e., we only need to look up few words tokens for a sentence in LLM; we only need to look up one user/item ID for a user-item training sample in RecSys), and ‘dense’ means the calculation flow parameters which will fully activate to estimate final results. Indeed, in terms of the amount of learnable parameters, our ranking model (¿200B) far exceeds many LLMs (around 100B), which indicates that learnable parameters are not the bottleneck of ranking models.
@@
  • For the computational complexity, FLOPs: Actually, RecSys model inference FLOPs is much lower than LLM and has upper bound. This is because our model needs to process hundreds of millions of requests efficiently, and that’s why our dense-activated DNN module is much smaller (about 0.1B). Thus, to ensure our service’s stability and robustness when processing each request in milliseconds (LLMs usually take seconds), we cannot blindly increase inference complexity. Additionally, since the real-time nature of training recommendation models greatly affects the quality of online performance, the computational complexity of model under limited offline training resources will impact the real-time nature of streaming training, resulting in trade-off with online performance, e.g, if the offline model has 20 minutes delay, it will cause 1% drop in online performance.
@@
Based on the analysis of the above three aspects, compared to LLM, we can observe that our ranking model has its advantages and disadvantages: (1) ‘unlimited’ streaming data, (2) massive parameter storage, (3) relatively low-FLOPs DNN module. In other words, for ranking model, data and storage resources are relatively cheap, but computing resources are cautiously expensive. Motivated by these points, we consider whether we can use the model’s strengths in data and storage to compensate for its weakness in computing. In other words, can we cache part of complex module calculation results to degenerate its time complexity?
[…]
@@
Fortunately, owing to the temporal order of streaming data widely used in industry RecSys, if we cache the frequently used masked SA module results on the latest samples, the complex masked SA layer can be replaced by simple TA layer to efficiently predict different item candidates well, as shown in Figure 2(d). In this way, our MARM extends the single-layer attention-based sequences interests modeling module to a multi-layer setting with a minor FLOPs cost. Specifically, our MARM solution significantly overcomes computational bottlenecks and could seamlessly empower all interest extraction modules for user sequences. Moreover, based on MARM cache results, we find it beneficial for other models, such as retrieval and cascading models. In our experiments, we fully explore scaling laws between cache size and model performance, seeking a perfect balance among cached attention depth, sequence length, and embedding dimension. Our main contributions are:
@@
  • We are the first work to explore new scaling-laws under the widely-used ranking model architecture from a fresh cache-performance perspective.
@@
  • We devise a simple-yet-efficient approach, MARM, which utilizes the strengths of ranking models in data and storage to reduce high time complexity of masked-self-attention. We successfully extend the single-layer attention-based sequences interests modeling module to a multi-layer setting with minor inference FLOPs.
@@
Figure 3. MARM realtime cache update&use workflow.
@@ 2. Methodology
In this section, we do not explain the total details of our ranking model’s architectures or losses. Specifically, for better understanding, we only dive into the MARM module and explain how it works, as sketched in Figure 2(d).
@@ 2.1. MARM Workflow
This section describes MARM workflow under the latest exposure item sequences, which builds a multi-layer decoder-only Transformer architecture to capture user interests with respect to a certain target item. Our MARM mainly consists of four parts:
@@
    An external cache memory storage to look up results.
@@
    Sending the intermediate results to the cache store.
@@ 2.1.1. Sequence Generator
where $n$ denotes sequence length, the $\texttt{Iid1},\dots,\texttt{Iidn}]$ is the behavior sequence for user UidX. Of course, since our sequence has clear user feedback, we can also choose to use items that meet certain conditions. For example, we might only use items that have a long-view label, see Section 2.2 for this extension.
@@ 2.1.2. Cache Memory Storage Look-Up
As shown in Figure 2, given UidX’s sequence $\texttt{Iid1},\dots,\texttt{Iidn}]$ and item candidate IidY, there are two types of representations needed to support our MARM: the learnable ID-based embedding and un-learnable cached results.
@@
    where $\mathbf{IidY}\in\mathbb{R}^{F}$, $F$ is a hyper-parameter to control the feature dimension. The $\mathbf{Iid1},\dots,\mathbf{Iidn}]$ is the the bottom target-attention input in Figure 2(a). It is worth noting that viewed items could fulfill some other attributes, such as author ID, tags, and user interaction labels. In this paper, for simplicity, we will only retain the item ID to represent them.
@@
    Un-Learnable Cached results: Considering we have cached $L$-layer intermediate results in an external key-value memory storage, we need to look up those results precisely. Therefore, we devise a hash strategy to generate ‘key’ to reflect the meaning: cached value of user-item pair at layer $i$. For example, given user UidX, and its item sequence [Iid1,…], to find cache results at layer depth $i$, we could generate hash key list as:
@@
    where the $\texttt{Iid1}^{i}_{\texttt{UidX}}\in\mathbb{Z}$ denotes the Hash-key to access the correct unique cache result. Based on them, we have:
@@
    where $\mathbf{Iid1}^{i}_{\texttt{UidX}}\in\mathbb{R}^{d}$, and $d$ is attention module dimension. As a result, combining all layers’ cache results, we have:
@@ 2.1.3. Multi-Layer Target-Attention
Up to now, we have obtained learnable target-attention (TA) ID embeddings and subsequent unlearnable TA cached results. According to learnable embeddings, we first calculate them with a certain target item information $\mathbf{IidY}$ through TA mechanism:
[…]
Next, we feed the results $\mathbf{IidY}_{\texttt{UidX}}^{1}\in\mathbb{R}^{d}$ to calculate with un-learnable cached results to simulate the high-FLOPs masked-self-attention (MSA) ($\mathcal{O}(n^{2}*d)$) by the low-FLOPs TA ($\mathcal{O}(n*d)$):
[…]
where $\mathbf{IidY}_{\texttt{UidX}}^{L+1}\in\mathbb{R}^{d}$ denotes final sequences interests modeling output belonging to (UidX, IidY) pair, and will be fed into following multi-task learning module to predict ground-truth labels.
2.1.4. Sending results to Cache Memory
The above sections have introduced our MARM calculating process. Next, we will explain how to save such intermediate results to the cache store. Similarly, we first generate the ‘hash key’ to indicate the result meanings, e.g., the $(\texttt{UidX},\texttt{IidY})$ result at different layers $1,\dots,L]$:
Afterwards, we can send the values with their keys in order:
In streaming training, as we continuously save real-time computation results, MARM cache gradually accumulates computation results of each layer of MARM for all users’ historical behaviors. Through natural accumulation of cache, MARM can apply decoder-only Transformer architecture with minor inference FLOPs. The pseudo-code of our MARM is shown in Algorithm 1, and Figure 3 also provides a user-watching example in streaming setting.
2.2. MARM with SIM
In Section 2.1, we have discussed a naive MARM setting, modeling the users’ interests under the latest watched items sequence. However, since the sequence length $n$ in Eq.(1) is limited (about 100), the earlier stored results will no longer be used once out of range, resulting in a waste of storage resources. This section further extends MARM with Search-based interests model (SIM)(Pi et al., 2020), to maximize the effectiveness of cached results.
Formally, SIM proposes a two-stage cascading modeling paradigm: (1) introduce a coarsen General Search Unit (GSU) to retrospect user life-long history (e.g. ¿10000) for target item, and then search the top-K most related item sequences; (2) employ fine-grained Exact Search Unit (ESU) to compress the searched sequence information to obtain precise user interests with respect to target item. In past years, SIM has been the vital engine for ranking model iteration, and the recent effort is TWIN (Chang et al., 2023), which utilizes the shared GSU and ESU modules for two-stage unbiased modeling. Inspired by the TWIN, we also consider applying a synchronous GSU module to search the top-K related cache results for each layer (as shown in Figure 4(a)). Given the latest long-term sequence $\texttt{Iid1},\dots,\texttt{Iid}10000]$ of user UidX, the MARM GSU aims to generate the multi-layer searched input sequences, for example:
[…]
where the MARMGSU shared same Target-Attention parameters with our ranking model, and $\texttt{S\_Iid1}^{i},\dots,\texttt{S\_IidK}^{i}]$ are top-K searched items with highest attention weights for candidate IidY at layer $i$. Combining all layers’ searched keys, we have:
$\footnotesize\begin{bmatrix}\texttt{S\_Iid1}^{1}&\texttt{S\_Iid2}^{1}&\texttt{S\_Iid3}^{1}&\dots&\texttt{S\_IidK}^{1}\\ \texttt{S\_Iid1}^{2}&\texttt{S\_Iid2}^{2}&\texttt{S\_Iid3}^{2}&\dots&\texttt{S\_IidK}^{2}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \texttt{S\_Iid1}^{L}&\texttt{S\_Iid2}^{L}&\texttt{S\_Iid3}^{L}&\dots&\texttt{S\_IidK}^{L}\end{bmatrix}$
where the searched sequence of different layer maybe different, e.g., $\texttt{S\_Iid}1^{i}\neq\texttt{S\_Iid}1^{i+1}$. Afterwards, we can look up their corresponding cached results to support our MARM module in Figure 4(b).
2.3. MARM to Support Other Models
Actually, the industrial RecSys always introduces several models to filter candidate items layer by layer:
  • Retrieval model: without item candidates, only utilizing users’ latest interaction items to fetch thousands of items from the full billion-item pool.
  • Cascading model: a small ranking model, utilizing extremely efficient modules to select top hundreds of item candidates from the thousands generated by the retrieval model.
  • Ranking model: in a broad sense, it is the most complex model, which finds the best dozens items from hundreds of candidates.
Generally, these models are trained using different strategies and are isolated from each other. Therefore, here is an interesting assumption: can our ranking model MARM cache storage contribute to other Retrieval/Cascading models? Unsurprisingly, we found our MARM results could support them seamlessly:
  • For Retrieval model, since there is no candidate set of items, we use users’ latest 200 watched items to look up the cache results from MARM, and then apply the user ID feature to replace item candidates to conduct target-attention as shown in Figure 4(c).
  • For Cascading model, we implement the same multi-layer target-attention process as our Ranking model in Eq.(6). It is important to note that in the cascading model, we do not perform re-accumulation of the MARM cache; instead, we directly use the already accumulated MARM cache and only use latest short-term sequences. This approach saves resources while also enhancing the consistency between cascading and ranking models.